Operations Report – November 13th

Crew 215 Operations Report 13-Nov-2019
SOL: 3
Name of person filing report: Shane Usher
Non-nominal systems: Opportunity, Curiosity
Notes on non-nominal systems: Suit 2 battery replaced. Opportunity broke down.
Curiosity plugged in with light flashing, but not charging battery.
Generator: run
Hours run: 14h 00 min
From what time last night: approx. 1745
To what time this morning: 0745
List any additional daytime hours when the generator was run: N/A
Solar— SOC % (Before generator is run at night): 72%
Diesel Reading – 50% –
Station Propane Reading – 81%.
Ethanol Free Gasoline: N/A
Water (loft tank): 47 gallons
Water Meter: 1457711 units
Water (static tank): 420
Static to Loft Pump used – yes
Water in Green Hab: 287.8 gallons, 20 L (5.7 gallons) used today
Water in Science Dome: 0 gallons
Toilet tank emptied: no
Deimos rover used: not at hab
Hours: –
Beginning charge: –
Ending charge: –
Currently charging: –
Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: –
Beginning charge: –
Ending charge: –
Currently charging: –
Spirit rover used: no
Hours: 110.5
Beginning charge: (Before EVA): 62%
Ending charge: (On return from EVA, before recharging): 100%
Currently charging: yes.
Opportunity rover used: yes
Hours: 66.2
Beginning charge: 100%
Ending charge: 55%
Currently charging: yes
Curiosity rover used: yes
Hours: 116.7
Beginning charge: (Before EVA) : 100%
Ending charge: (On return from EVA, before charging): 67%
Currently charging: yes
Notes on rovers: Opportunity broke down. Curiosity does not appear to be charging,
light is flashing but % charge is not increasing.
ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): no
Reason for use: N/A
Oil Added? no
ATV Fuel Used: 0 Gals
# Hours the ATVs were used today: 0
Notes on ATVs: Assigned to the director
HabCar used and why, where? With DIRECTOR
CrewCar used and why, where? At Grand Junction awaiting Crew 216
General notes and comments: N/A
Summary of internet: Functional, but there was a period of no internet access.
Summary of suits and radios: Suit 2 battery replaced, Suit 5 Helmet photo sent.
Summary of Hab operations: No new issues.
Summary of GreenHab operations: Atmospheric equalization to the green hab was carried
out during the day, plants were watered (20L). The heater automatically came on at 5pm.
Summary of Science Dome operations: All systems nominal.
Summary of RAM operations: Nominal
Summary of any observatory issues: Nominal
Summary of health and safety issues: All crew healthy.
Questions, concerns and requests to Mission Support: Opportunity now may not be usable.
Waiting to see if Curiosity is charging.

Copyright © The Mars Society. All rights reserved. | Main Site